# Fairness-in-the-Workflow: How Machine Learning Practitioners at Big Tech Companies Approach Fairness in Recommender Systems

### v2

@@ 7. Calls to Action
In this section, we map the pain points uncovered in our interviews to calls for action, targeted towards RS and fairness communities alike.
7.1. Building Institutional Knowledge Through Documentation and Collaboration
@@
A key challenge reported by RS practitioners was adapting fairness metrics that had been previously developed at their organizations for new fairness work. Teams are composed of individuals from a wide variety of personal backgrounds, numerous project stages are spread over time, and systems are improved dynamically after deployment, when fairness effects may be hard to predict and can potentially self-reinforce in feedback loops. Due to these challenges, documenting prior fairness efforts is critical for building institutional fairness knowledge. However, as multiple practitioners pointed out, while there are standard practices for documenting and approving code changes, there is comparably little infrastructural support for documenting approaches to assess and improve fairness. Researchers have previously called for such fairness documentation (Heger et al., 2022), e.g., in the style of datasheets for datasets (Gebru et al., 2021; Papakyriakopoulos et al., 2023). An internal template for fairness documentation for RS can help to improve the process in the immediate term and also build institutional knowledge.
Considering that RS involve massive amounts of content, large user bases, rapidly changing environments, and large teams of developers, it is surprising that the practitioners we interviewed did not identify stronger support for maintaining team cohesion and situational awareness during project development. While collaboration was explicitly cited as a stage of the fairness process by practitioners, practitioners did not report receiving specific technical or procedural support to facilitate cross-team collaboration. We believe there is a need for research into best practices for collaboration and information exchange during RS development to better support cross-team fairness efforts.
7.2. Encouraging Researcher and Organizational Investment in Multi-Stakeholder Fairness
@@
To tackle the challenges involved in improving fairness in RS, organizations must invest intellectual resources into researching, or working to implement academic research on, several key areas. Firstly, it is necessary to develop more comprehensive optimization approaches, considering multiple objectives, to balance (or avoid) tradeoffs between different stakeholder groups. Secondly, further investigations into group-wise fairness performance comparisons should address the “too many groups” challenge in RS by making it easier to assess fairness across a very high numbers of groups. Similarly, there should also be an increased focus on intra-group fairness to align with multi-faceted and intersectional conceptions of diversity in the real world. Additionally, researchers and organizations should work to develop standardized guidelines for making tradeoff decisions to scaffold decision-making when fairness tradeoffs between stakeholder groups is unavoidable. This would help avoid the ad hoc and potentially arbitrary decision-making that currently occurs when RS practitioners are forced to make these decisions in real time as tradeoffs arise. Guidelines should be developed in collaboration with diverse stakeholder groups, and should encourage practitioners to question, rather than reinforce, existing power dynamics on RS platforms.
@@ 7.3. Integrating Fairness Considerations into Workflows Early and Often
On an organizational level, we found that practitioners encountered competing incentives for integrating fairness into their workflows. While practitioners’ primary job responsibilities were to create and manage RS, every practitioner interviewed felt both a personal interest and an explicit requirement to assess and improve fairness in their models. At the same time, however, we noticed many cases where practitioners had to backtrack or iterate on deployed models in order to mitigate unexpected instances of unfairness. Had they invested additional time during project formation and received support from fairness teams earlier, perhaps such issues could have been detected when changes would have been easier to make. Our workflow models (Figures 1 and 2) highlight the opportunity for early engagement with fairness teams, at transition points in the workflow and during phases of the RS workflow that align with phases of the fairness process. We also call for additional research into the organizational context of RS in order to identify organizational barriers beyond the scope of our investigation.
@@ 7.4. Developing a ‘Fairness Lingua Franca’
Across Industry Teams
@@
We identified cases where poor communication between fairness teams and technical practitioners led to confusion and even deprioritization of fairness efforts. While legal teams offered explicit guidance and directives to practitioners, the nature of fairness feedback—reliant on nuance and not necessarily immediately applicable to complex models—proved much harder to parse and operationalize in systems. Combined with the fast-paced nature of RS development, this led to disconnects between technical practitioners and fairness teams despite mutual agreement on the importance of fairness. We believe that these disconnects could have been anticipated. Practitioners reported that fairness teams came from multiple disciplines, including some (like psychology and communication) which may not share common terminology with ML. RS practitioners lacked the broad interdisciplinary expertise represented on fairness teams, resulting in a disconnect in fully understanding the implications and context of fairness feedback. The dynamic environment of RS exacerbated this gap, leading practitioners to optimize by solving easy issues while neglecting other—potentially more serious—issues. We found that the academic literature was often viewed as a reliable, albeit not universally applicable, resource by both practitioners and fairness teams. We therefore posit that HCI and ML researchers might also help with the necessary ‘translation’ work to provide useful vocabulary and common ground for practitioners. Likewise, embedding fairness practices more deeply in technical teams through additional roles, building domain expertise in mutual fields, and pursuing techniques for more consistent documentation and information flow (especially at the beginning of the RS workflow) might also help to establish common ground.
Across Industry and Academia
We found that knowledge from research communities not only shaped many practitioners’ awareness of fairness but also gave them some basic principles to follow. Yet practitioners also regularly developed their own ad hoc fairness metrics instead of using those described in academic literature. This creates a patchwork of practices that at times adheres to academic recommendations and at other times drifts far from them. While practitioners do not necessarily seek to disregard academic best practices, practical constraints sometimes make it hard to follow, or even known about, them. We call for greater collaboration between industry and academia, acknowledging that accomplishing this goal is challenging and limited by many factors beyond the scope of our study. However, there are clear benefits to both sides: industry practitioners may benefit from investigating novel approaches, while research communities can potentially gain exposure to new use cases and novel empirical evidence.
@@ 8. Discussion
8.1. The Path Forward: RS Practitioners Emphasize Benefits of Fairness Work Despite Challenges
Despite the aforementioned challenges, we have reason to be optimistic: RS practitioners consistently ascribe importance to fairness work. In our semi-structured interviews, we asked practitioners to describe their emotional stress and satisfaction level in their daily work. While most practitioners reported that balancing daily work and fairness considerations was somewhat stressful, most also believed that fairness is an important and necessary part of their work. P9 described incorporating fairness as “…tiring but really rewarding.” Practitioners pointed out that fairness felt particularly important given the uniquely large impact that RS play on society. P1 notes: “It is tiring; however, I believe this is the right thing to do. It is important to remove bias as much as possible as our system serves millions of users worldwide. Discriminating behaviors imposed by our system might greatly affect their life.” Practitioners expressed a desire to improve the world through their systems and identified strong intrinsic motivations for improving fairness, even if it is arduous. Some practitioners cited public and academic discourse on fairness as motivating; for instance, P1 mentioned, “Thorough discussions from existing papers, you know, like COMPAS, about how unfairness affects other people’s life greatly shaped my view on this. As I have the chance to repair this, I will do my best—though it is not easy.”
@@ 8.2. Limitations and Ethical Considerations
Our study is subject to a number of limitations. First and foremost, our findings and discussions are based on a limited sample of practitioners from seven highly visible companies. Although our sample size is typical of HCI interview studies (Caine, 2016), and although we intentionally targeted practitioners working on very large-scale RS with high societal impact, this limits the kind of evidence that we were able to gather. In particular, our findings may not generalize to RS practitioners working on smaller-scale RS. Similarly, this paper focuses on machine learning practitioners working on RS who are leading engineering or research efforts; we did not interview members of legal, data, or fairness teams. The desires we uncovered may also not be fully actionable: there is a fundamental tension between the desire for fairness rules that function similarly to rigid and straightforward legal and data governance requirements, and the ability actually to implement such rules in the wild (Ali et al., 2023). Finally, our semi-structured interviews might be limited by the remote setting, potential risks of exposing corporate secrets, or participant self-presentation maintenance. While we found participants to generally be quite frank about their experiences, it is possible that they provided a more critical or a rosier picture of the environment due to their perceived expectations.
@@
While this study aims to foster understanding and improve fairness in RS, we acknowledge the potential for adverse impacts once published. The findings could unintentionally influence RS practitioners to adopt fairness practices that may not be universally applicable, leading to oversimplification of complex issues. Furthermore, there is a risk that the emphasis on certain strategies might overshadow other equally important aspects of fairness in RS. We recognize the importance of ongoing critical evaluation and discourse in the community to mitigate these potential negative impacts and encourage a diverse range of perspectives and solutions in addressing structural issues of fairness in recommender systems.
@@ 9. Conclusion
In this work, we investigated fairness practices in RS through a semi-structured interview study of 11 technical practitioners who are building recommender systems. We constructed a process-level map of RS practitioners’ daily workflow, focusing on how they interact with legal, data, and fairness teams at their organizations and on how they incorporate fairness into their existing practices. We identified key technical challenges to incorporating fairness, such as defining fairness in the context of RS and navigating the complexity introduced by RS; namely, multi-stakeholder fairness considerations and dynamic environments. We also identified organizational challenges resulting from cross-team interactions, including a lack of time to conduct fairness work and an communication challenges across teams stemming from a lack of a ‘fairness lingua franca.’ We suggest that building institutional knowledge by improving documentation and collaboration practices, encouraging researcher and organizational investment into multi-stakeholder fairness approaches, integrating fairness considerations into RS workflows early and often, and developing a ‘fairness lingua franca’ can lead to improvements in RS fairness.
@@ Acknowledgements.
This work was supported by NSF grant IIS-1850195. We would like to thank the associate chairs and anonymous reviewers for their invaluable feedback. We also would like to thank Dataprep team from Simon Fraser University for the valuable input of mapping of common functions and user intentions. All opinions, findings, and conclusions in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.

### v3

@@ 7. Calls to Action
In this section, we map the challenges uncovered in our interviews to calls for action, highlighting both where big tech companies and practitioners are responsible for implementing evidence-based changes to better integrate fairness into their RS workflows as well as where novel HCI research can support real-world fairness efforts.
7.1. Building Institutional Knowledge Through Documentation
@@
A key challenge reported by RS practitioners was adapting fairness metrics that had been previously developed at their organizations for new fairness work. Teams are composed of individuals from a wide variety of personal backgrounds, numerous project stages are spread over time, and systems are modified dynamically after deployment, when fairness effects may be hard to predict and can potentially self-reinforce in feedback loops. Due to these challenges, documenting fairness work is critical for building institutional fairness knowledge. However, as multiple practitioners pointed out, while there are standard practices for documenting and approving code changes, there is comparably little infrastructural support for documenting approaches to assess and improve fairness.
The Role of Practitioners and Organizations
Practitioners and organizations must document their fairness metrics and prior fairness work. Practitioners and organizations are also responsible for effectively building on institutional knowledge. For example, the practitioners we interviewed reported that, when available, they re-used metrics from prior projects in the initial prototyping stage of the fairness workflow, but typically developed around five brand-new metrics by the end of the collaboration stage. This ad-hoc approach impedes the accumulation of institutional knowledge. It may harm procedural fairness as well, as different fairness standards are applied across different systems, and standards may not always be rigorously evaluated for validity. Thus, we suggest that practitioners should design metrics with extensibility, or the ability for metrics to be “adapted for different systems, use cases, and deployment contexts” in mind (Harvey et al., 2025).
The Role of HCI Researchers
HCI and ML fairness researchers have previously led calls for improved documentation in both research and industry. Researchers have proposed documentation frameworks including datasheets for datasets (Gebru et al., 2021) and model cards for model reporting (Mitchell et al., 2019)—and have led efforts to understand the extent to which documentation meets the needs of and is adopted by industry practitioners (Heger et al., 2022; Boyd, 2021). More recently, researchers have begun to propose domain-specific augmentations to documentation frameworks, such as datasheets for speech datasets (Papakyriakopoulos et al., 2023) or for digital cultural heritage datasets (Alkemade et al., 2023). Others have documented how AI practitioners themselves customize and adapt artifacts from research, like fairness checklists, to suit their needs and align with their workflows (Madaio et al., 2024; Morley et al., 2023; Madaio et al., 2020). Going forward, HCI researchers are uniquely well-suited to determining how such artifacts should be designed, i.e., which design practices lead to the development of adaptable or extensible documentation frameworks. Some researchers are already working on this: for example, Madaio et al. (2024) suggest that designers of artifacts intended for use by industry practitioners embrace “positive ambiguity,” or under-specification, so that artifacts can more easily be adapted. Similarly, Harvey et al. (2025) propose that artifacts should be designed to be modifiable and modular in order to achieve extensibility. Both Madaio et al. (2024) and Harvey et al. (2025) leave open the question of how ambiguous or modifiable documentation frameworks and other artifacts can be before they no longer perform their intended function of standardization, suggesting one clear avenue for future HCI research.
7.2. Investing in Multi-Stakeholder Fairness
Multi-stakeholder fairness considerations raised a variety of challenges, including managing conflicting stakeholder interests; prioritizing between many different, non-conflicting, types of fairness; and managing the technical and prioritization challenges associated with handling high numbers of user groups.
The Role of Practitioners and Organizations
@@ 7.2. Investing in Multi-Stakeholder Fairness
To tackle the challenges involved in improving fairness in RS, organizations must invest resources into researching, or working to implement academic research on, several key areas. First, it is necessary to develop more comprehensive optimization approaches, considering multiple objectives, to balance (or avoid) tradeoffs between different stakeholder groups. Second, research should address the “too many groups” challenge in RS by making it easier to assess fairness across a very high numbers of groups. Similarly, there should be an increased focus on intra-group fairness to align with multi-faceted and intersectional conceptions of diversity in the real world (Wang, 2025; Wang et al., 2022; Kanubala and Valera, 2025). While ML researchers can support this research, the massive scale of modern RS mean that much of this work will be computationally and financially expensive, and thus much of the responsibility should fall to well-resourced big tech organizations who also profit from RS (Widder et al., 2024).
The Role of HCI Researchers
HCI researchers can advance multi-stakeholder fairness by developing transparent and standardized guidelines to scaffold decision-making when fairness tradeoffs between stakeholder groups are unavoidable. Value-sensitive design (VSD) methods, which embrace tensions between conflicting stakeholder values as a core theoretical construct, could prove a valuable approach (Friedman and Hendry, 2019; Boyd, 2022; Zhu et al., 2018; Simon et al., 2020; Shen et al., 2022). Both Zhu et al. (2018) and Simon et al. (2020) have suggested that VSD’s methodology of iterative conceptual, empirical, and technical investigations can be incorporated into the design of fair algorithms. As a concrete example, Shen et al. (2022) use a VSD approach to create the Model Card Authoring Toolkit, which is intended to help community members with “tradeoff deliberation” while designing decision-making algorithms. We suggest that HCI research focused on similar VSD-informed efforts could help avoid the ad hoc and potentially arbitrary decision-making that currently occurs when RS practitioners are forced to make fairness tradeoffs in real time. Guidelines should be developed in collaboration with diverse stakeholder groups, and should encourage practitioners to question, rather than reinforce, existing power dynamics on RS platforms.
@@ 7.3. Integrating Fairness Considerations into Workflows Early and Often
RS practitioners typically did not interact with fairness teams until models had been deployed. This delay caused friction, as fairness teams sometimes had out-of-date information on technical terminology or system updates. At the same time, during online phases of model development (when fairness teams were in the loop), technical teams were consistently “on-call” (P2) and reported prioritizing maintaining model uptime over fairness fixes.
The Role of Practitioners and Organizations
On an organizational level, we found that practitioners encountered competing incentives for integrating fairness into their workflows. While practitioners’ primary job responsibilities are to create and manage RS, every practitioner interviewed felt both a personal interest and an explicit requirement to assess and improve fairness in their models. At the same time, however, we noticed many cases where practitioners had to backtrack or iterate on deployed models in order to mitigate unexpected instances of unfairness. Had they invested additional time during project formation and received support from fairness teams earlier, perhaps such issues could have been detected when changes were easier to make. Our workflow model (Figures 1) highlights the opportunity for earlier engagement with fairness teams. The practitioners we interviewed explicitly stated that they would have preferred fairness teams to be involved with their projects earlier on; there is ample opportunity for this during the prototyping and internal phases of the fairness workflow.
The Role of HCI Researchers
Researchers are well-equipped to develop tools that support industry practitioners in incorporating fairness into their workflows. For example, Boenig-Liptsin et al. (2022) have proposed the Ethos Lifecycle, an interactive online tool designed to increase the extent to which data science practitioners consider positionality, power, the interrelatedness of social and technical considerations, and the narratives animating their work. The tool is built around a model of a data science workflow and invites users to consider relevant ethical questions at each stage of the workflow. Similarly, Boyd (2022) uses VSD principles to create a field guide to inform practitioners about technical mitigations to ethical issues that they identify in the course of their work. That research was grounded in the concept of ethical sensitivity, which Boyd and Shilton (2021) define as “the moment of noticing an ethical problem (recognition), the process of building understanding of the situation (particularization), and the decision about what to do (judgment).” Following Boyd and Shilton (2021), we suggest that additional HCI research is needed to understand the ways in which ethical recognition can be triggered and how such triggers can be built in to industry workflows. Furthermore, we suggest that HCI researchers can build on our model of the RS practitioner workflow to create teaching tools that could improve ethical particularization and judgment and encourage early consideration of fairness in the development of RS.
@@ 7.4. Developing a ‘Fairness Lingua Franca’
We identified several cases where poor communication between fairness teams and technical practitioners led to confusion and even deprioritization of fairness efforts.
The Role of Practitioners and Organizations
@@
While legal teams offered explicit guidance and directives to practitioners, the nature of fairness feedback—reliant on nuance and not necessarily immediately applicable to complex models—proved much harder to parse and operationalize in systems. Combined with the fast-paced nature of RS development, this led to disconnects between technical practitioners and fairness teams despite mutual agreement on the importance of fairness. We believe that these disconnects could have been anticipated. Practitioners reported that fairness teams came from multiple disciplines, including some (like psychology and communication) which may not share common terminology with ML. RS practitioners lacked the broad interdisciplinary expertise represented on fairness teams, resulting in barriers to fully understanding the implications and context of fairness feedback. The dynamic environment of RS exacerbated this gap, leading practitioners to optimize by solving easy issues while neglecting other—potentially more serious—ones. Nevertheless, the gap could be reduced by more frequent meetings between technical and fairness teams in which shared terminologies are established. Likewise, embedding fairness practices more deeply in technical teams through additional roles, building domain expertise in mutual fields, and pursuing techniques for more consistent documentation and information flow (especially at the beginning of the RS workflow) might help to establish common ground.
We found that, when participants were aware of RS fairness research, research could provide basic principles for practitioners to follow or represent starting points for bespoke metrics. Yet several practitioners reported that they did not have the time to regularly review academic research, and thus that they consistently developed their own ad hoc fairness metrics instead. This creates a patchwork of practices that at times adheres to academic recommendations and at other times drifts far from them. While practitioners do not necessarily seek to disregard academic best practices, practical constraints sometimes make it hard to follow, or even know about, them. We suggest that the big tech companies that employ RS practitioners are responsible for incentivizing practitioners to devote the time necessary to stay up-to-date on academic research, i.e., by making it an explicit job responsibility as opposed to a form of invisible labor (Deng et al., 2023b).
The Role of HCI Researchers
We suggest that HCI researchers might help with the necessary translation work to provide common ground for technical and fairness practitioners. For example, Morley et al. (2021) propose Ethics as a Service, as a model for helping practitioners “translate principles into practice.” Ethics as a Service suggests concrete roles for organizations and practitioners to adopt organization-wide ethical principles and determine the appropriate approaches for implementing them. Similarly, Delgado et al. (2022) identify factors that led to the successful translation of real-world legal requirements into an AI system through a participatory design task, including extended co-design and evaluation of the AI solutions and the creation of opportunities for expertise to “flow in multiple directions.” HCI researchers can build on this prior work to identify best practices for collaboration and information exchange during RS development to better support cross-team fairness efforts. A particularly useful starting point for this work is Deng et al. (2023b), who describe how practitioners take on “bridging” roles to build AI fairness expertise at their organizations and identify a variety of efforts undertaken by practitioners with fairness expertise to educate practitioners without that expertise, including hosting design workshops and hackathons and developing guidebooks and toolkits. At the same time, the work finds that bridging efforts are labor-intensive and under-appreciated (Deng et al., 2023b). Thus, we call for additional research into the effectiveness of different bridging strategies, as well as how bridging work can be made easier for practitioners.
@@ 8. Discussion
Despite the aforementioned challenges, we have reason to be optimistic: RS practitioners consistently ascribe importance to fairness work. In our semi-structured interviews, we asked practitioners to describe their emotional stress and satisfaction level in their daily work. While most practitioners reported that balancing daily work and fairness considerations was somewhat stressful, most also believed that fairness is an important and necessary part of their work. P9 described incorporating fairness as “tiring but really rewarding.” Practitioners pointed out that fairness felt particularly important given the uniquely large impact that RS play on society:
P1: It is tiring; however, I believe this is the right thing to do. It is important to remove bias as much as possible as our system serves millions of users worldwide. Discriminating behaviors imposed by our system might greatly affect their lives.
Practitioners expressed a desire to improve the world through their systems and identified strong intrinsic motivations for improving fairness, even if it is arduous. Some practitioners cited public and academic discourse on fairness as motivating:
P1: Thorough discussions from existing papers, like COMPAS (Angwin et al., 2016), about how unfairness affects other people’s life greatly shaped my view on this. As I have the chance to repair this, I will do my best—though it is not easy.
@@ 8.1. Limitations and Ethical Considerations
Our study is subject to a number of limitations. First and foremost, our findings and discussions are based on a limited sample of practitioners from seven highly visible companies. Although our sample size is typical of HCI interview studies (Caine, 2016), and although we intentionally targeted practitioners working on very large-scale RS with high societal impact, this limits the kind of evidence that we were able to gather. In particular, our findings may not generalize to RS practitioners working on smaller-scale RS. Similarly, it is likely that fairness is more salient to the specific subset of individuals who agreed to be interviewed than it would be to a representative sample of RS practitioners. This does not necessarily imply that the individuals we interviewed feel differently about fairness than the average RS practitioner, but it is certainly possible that they were more likely to believe that fairness was “rewarding” (P9) and “the right thing to do” (P1) (see §8). In any case, we do not attempt to draw generalizations about how common the fairness challenges and beliefs we identify are among RS practitioners as a whole—we only claim that these challenges exist, and that they impact the development of large-scale RS. Furthermore, this paper focuses on technical practitioners working on RS who are leading engineering or research efforts; we did not interview members of legal, data, or fairness teams. We chose to do this because we were specifically interested in how practitioners who are actively involved in building RS incorporate fairness into their responsibilities. Nevertheless, this does prevent us from analyzing the extent to which the workflows described by legal, data, and fairness teams align with—and deviate from—those described by technical teams. Further, our semi-structured interviews might be limited by the remote setting, potential risks of exposing corporate secrets, or participant self-presentation maintenance. While we found participants to generally be quite frank about their experiences, it is possible that they provided a more critical or a rosier picture of the environment due to their perceived expectations. Finally, the challenges we uncovered are likely not fully addressable: there is a fundamental tension between the desire for fairness rules that function similarly to rigid and straightforward legal and data governance requirements, and the ability to actually implement such rules in the wild (Ali et al., 2023).
@@
While this study aims to foster understanding and improve fairness in RS, we acknowledge the potential for adverse impacts once published. The findings could unintentionally influence RS practitioners to adopt fairness practices that may not be universally applicable, leading to oversimplification of complex issues. Furthermore, there is a risk that the emphasis on certain strategies might overshadow other equally important aspects of fairness in RS. We recognize the importance of ongoing critical evaluation and discourse in the community to mitigate these potential negative impacts and encourage a diverse range of perspectives and solutions when addressing structural issues of fairness in RS.
@@ 9. Conclusion
In this work, we investigated fairness practices in RS through a semi-structured interview study of 11 technical practitioners who are building RS. We constructed a map of RS practitioners’ daily workflow, focusing on how they interact with legal, data, and fairness teams at their organizations and on how they incorporate fairness into their existing practices. We identified key technical challenges to incorporating fairness, such as defining fairness in the context of RS and navigating the complexity introduced by RS; namely, multi-stakeholder fairness considerations and dynamic environments. We also identified organizational challenges resulting from cross-team interactions, including a lack of time to conduct fairness work and communication challenges across teams. We suggest that building institutional knowledge by improving documentation practices, encouraging investment into multi-stakeholder fairness approaches, integrating fairness considerations into RS workflows early and often, and developing a ‘fairness lingua franca’ can lead to improvements in RS fairness.
@@ Acknowledgements.
This work was supported by NSF grant IIS-1850195. We would like to thank the associate chairs and anonymous reviewers for their invaluable feedback. We also would like to thank the Dataprep team from Simon Fraser University for the valuable input of mapping common functions and user intentions. All opinions, findings, and conclusions in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.
Appendix A Semi-Structured Interview Guide
Interview Phase Question Type Questions
Ice-Breaking Participant background Please tell me a bit more about yourself.
Participant company Please tell me something about the recommender system that you are helping to build.
Participant experience How long have you been working in this field?
How long you have been working at your current company?
Are you focusing on research, production, or a mixture of both?
Workflow and Approach to Fairness Participant Workflow (opening questions) Could you describe a recent project’s workflow?
Are there external teams that you interact with regularly?
Could you categorize your basic daily stages?
How do you feel about your workflow?
Participant Workflow (follow-up questions) Which stage do you think you spend most of your time on?
Would you mind sharing a recent interaction with [X] team?
How do you feel while interacting with [X] team?
Participant Approach to Fairness How do you define bias in your daily work?
Where do you find metrics and tools to quantify and detect bias?
Can you share a recent experience detecting bias in the RS that you work with?
Do you rely on a taxonomy of unfairness or harm in your work? If yes, how do you find it useful?
What amount of time do you spend on fairness in your daily work?
How do you feel about your experience with fairness in RS?
Organizational Approach to Fairness Is fairness in your organization a team effort, or is responsibility for fairness distributed just to one person or a small group? What do you think of this assignment?
In terms of mitigation approaches, do you prefer model-centric ones or data-centric ones? Why?
Could you describe a recent time that improving fairness failed or was extremely difficult?
What are the barriers you find that will prevent you from improving fairness?
How do you like the current partition of the time? Would you prefer to spend more or less effort on fairness? Why?
Work Experience Participant Preferences Describe fairness-related tools that you are familiar with. Which tools are you more or less likely to use? Why?
Can you rate your professional experiences in RS in terms of emotional support, work intensity, and stress level?
Table 3. Semi-Structured Interview Guide. As is typical of semi-structured interview studies, not all participants were asked the exact same set of questions in the same order. Some participants were asked additional follow-up questions as appropriate.
A semi-structured interview guide. It includes: (1) ice-breaking questions: Please tell me a bit more about yourself, Please tell me something about the recommender system that you are helping to build, How long have you been working in this field?, How long you have been working at your current company?, Are you focusing on research, production, or a mixture of both? (2) Questions about participants’ workflows and approaches to fairness: Could you describe a recent project’s workflow?, Are there external teams that you interact with regularly?, Could you categorize your basic daily stages?, How do you feel about your workflow?, Which stage do you think you spend most of your time on?, Would you mind sharing a recent interaction with [X] team?, How do you feel while interacting with [X] team?, How do you define bias in your daily work?, Where do you find metrics and tools to quantify and detect bias?, Can you share a recent experience detecting bias in the RS that you work with?, Do you rely on a taxonomy of unfairness or harm in your work? If yes, how do you find it useful?, What amount of time do you spend on fairness in your daily work?, How do you feel about your experience with fairness in RS?, Is fairness in your organization a team effort, or is responsibility for fairness distributed just to one person or a small group? What do you think of this assignment?, In terms of mitigation approaches, do you prefer model-centric ones or data-centric ones? Why?, Could you describe a recent time that improving fairness failed or was extremely difficult?, What are the barriers you find that will prevent you from improving fairness?, How do you like the current partition of the time? Would you prefer to spend more or less effort on fairness? Why? (3) Work experience questions: Describe fairness-related tools that you are familiar with. Which tools are you more or less likely to use? Why? Can you rate your professional experiences in RS in terms of emotional support, work intensity, and stress level?